OpenStudy (anonymous):

write (1-root3i)^8 in the standard form a+bi

OpenStudy (anonymous):

There's a couple ways to do this - have you learned binomial expansion?

OpenStudy (anonymous):

well, if not, you can muscle your way through it \[(1- \sqrt{3}i)^8 = \big ((1- \sqrt{3}i)^2 \big)^4\] \[ = (1-2 \sqrt{3}i-3)^4 \\ \ \\ =(-2-2\sqrt{3}i)^4 \\ \ \\ =(-2)^4(1+ \sqrt{3}i)^4 \\ \ \\ =16\big( (1+ \sqrt{3}i)^2\big)^2 \\ \ \\ =16\big( (1+2 \sqrt{3}i-3)\big)^2 \\ \ \\ = .......\] Can you finish it from there?

OpenStudy (anonymous):

remember that \[\sqrt{-1} =i \\ \ \\ i^2=-1\]